# Fast and efficient formation of stable tetraatomic molecules from ultracold atoms via generalized stimulated Raman exact passage

### v1

@@ Abstract
The study of the conversion of ultracold atoms into molecules has long remained a hot topic in atomic, molecular, and optical physics. However, most prior research has focused on diatomic molecules, with relatively scarce exploration of polyatomic molecules. Here we propose a two-step strategy for the formation of stable ultracold tetraatomic molecules. We first suggest a generalized nonlinear stimulated Raman exact passage (STIREP) technique for the coherent conversion of ultracold atoms to tetraatomic molecules, which is subsequently followed by a chainwise-STIREP technique to transfer the resulting molecules into a sufficiently stable ground state. Through systematic numerical analysis, we demonstrate that the proposed two-step strategy holds great potential for enabling the robust, fast, and efficient formation of stable ultracold tetraatomic molecules.
@@ I Introduction
The successful creation of various ultracold diatomic molecules has stimulated growing interest in polyatomic molecules (UPMs) Yang et al. (2022); Vilas et al. (2024); Chen et al. (2024). In contrast to diatomic molecules, UPMs offer additional complexity Carr et al. (2009); Küpper et al. (2009); Patterson et al. (2010), lending themselves to numerous interesting studies in, for instance, cold chemistry Karman et al. (2024), precision measurements Hutzler (2020) and the realization of exotic quantum phases Schmidt et al. (2022), to name a few. In order to tap their full potential, however, long-lived ensembles and full quantum state control serve as prerequisites Langen et al. (2024). Thus, the efficient conversion of ultracold atoms into stable UPMs constitutes an essential requirement, which hitherto remains challenging.
@@
In the ongoing experiments, ultracold molecules are first prepared by association of pre-cooled atoms Chin et al. (2010), either via photo- or the magneto-association Dulieu and Gabbanini (2009); Chandler (2010); Quéméner and Julienne (2012); Zhao and Pan (2022). Subsequently, the formed molecules should be immediately transferred to absolute ground states, it is commonly achieved via the well-known Stimulated Raman adiabatic passage (STIRAP) and its variants, such as multi-state STIRAP Danzl et al. (2010); Kuznetsova et al. (2008); Jing et al. (2008); Qian et al. (2010); Devolder et al. (2021). In principle, by manipulating the system to evolve along the adiabatic dark state, an efficient and robust state transfer can be obtained Bergmann et al. (1998); Král et al. (2007); Vitanov et al. (2017). Until now, STIRAP has been demonstrated successfully Bergmann et al. (2019), however, its application has been limited to the realm of diatomic molecules Winkler et al. (2007); Lang et al. (2008); Danzl et al. (2008); Aikawa et al. (2010); Ospelkaus et al. (2008); Feng et al. (2009); Liu et al. (2019a); Duda et al. (2023); Zamarski et al. (2025). Directly applying it to UPMs preparation remains challenging. To address this problem, researchers have developed STIRAP variants, for instance, Feshbach/Efimov resonance-assisted STIRAP, to support theoretical investigations into UPMs preparation Jing et al. (2007); Meng et al. (2009); Li and Peng (2011); Dou et al. (2012). Nevertheless, the successful application of such techniques is limited by the adiabatic criterion, which renders the systems vulnerable to environmental factors, e.g., collisions or decay to outside states.
@@
In order to relax limitations such as adiabaticity related to the STIRAP-based scheme, a technique which is called shortcut-to-adiabaticity (STA) was proposed Chen et al. (2010). By leveraging the STA, the adiabatic process can be expedited while ensuring that the outcomes remain consistent. One particularly successful method for designing such fast routes is counterdiabatic driving Demirplak and Rice (2005, 2008); Berry (2009), which allows one to construct a modification of an original Hamiltonian to eliminate unwanted nonadiabatic couplings Guéry-Odelin et al. (2019). Currently, STA is widely used in theory to realize fast and robust ultracold molecule formation Zhu and Chen (2021); Zhang and Dou (2021); Cao et al. (2023) and control the selective vibrational population transfer of molecules Masuda and Rice (2015a, b). In a recent study, Zhu etet alalZhu and Chen (2021) proposed a fast-forward scaling-assisted STIRAP method to study the conversion from atoms to diatomic molecules. In another study, Vitanov developed a shortcut multi-state STIRAP technique Vitanov (2020), which may provide an alternative ways for forming stable ground-state ultracold molecules. Compared to the traditional three-level system, the employment of multi-level ones exhibits unique advantages; to illustrate, this strategy proves instrumental in resolving the challenge of weak Franck-Condon factors (FCFs) during the process of both atom-to-molecule conversion Jaksch et al. (2002); Qian et al. (2010) and ground-state molecular transfer Danzl et al. (2010); Kuznetsova et al. (2008). Despite these advances, the major problem is that most prior studies usually require introducing additional couplings for eliminating non-adiabatic effects, which may pose additional experimental challenges. On the other hand, in 2017, Dorier etet alalDorier et al. (2017) suggested an alternative method, namely nonlinear stimulated Raman exact passage (STIREP), for the efficient atom-to-molecule conversion. This method conforms to the spirit of STA, because the control fields are determined by the specific dynamics selected to achieve exact state transfer Laforgue et al. (2022). This trick has been extended to the formation of triatomic molecules Cao (2024). Currently, an important question in this realm is whether it is possible to extend this successful tool to form larger polyatomic molecular complexes Dou et al. (2013); Gacesa et al. (2021); Shammout et al. (2025), and generalize it to transfer of molecules into deeply-bound states Danzl et al. (2010); Kuznetsova et al. (2008); de Lima (2024); Mukherjee and Tomza (2025).
@@
The purpose of this paper is to develop a theoretical formalism that extends the existing STIREP protocols to the case that permits control of conversion from atoms to stable tetraatomic molecules. We first report a generalization of the STIREP technique for controlling the conversion from ultracold atoms to tetraatomic molecules. To transfer the resulting molecules to a sufficiently stable ground state in a short time, we further develop a chainwise-STIREP (abbreviated as C-STIREP hereafter) technique. We present the complete explanations of the underlying mechanism and the corresponding numerical analyses. The results demonstrate that our proposed two-step strategy performs well in both the preparation of ultracold tetraatomic molecules and the ground-state molecular transfer.
@@
II COHERENT CONVERSION FROM ATOMS TO TETRAATOMIC MOLECULES VIA GENERALIZED NONLINEAR STIREP
@@ II.1 Model and Method
As outlined in the introduction, our top priority is to prepare homonuclear tetraatomic molecules. Figure. 1(a) shows a visualization of the process.
Figure 1: (Color online) Schematic showing the coherent conversion from ultracold atoms to tetraatomic molecules.
Starting from a Bose condensate of atoms, we apply the field $\Omega_{1}^{\prime}$ to associate free atoms $|a\rangle$ into triatomic molecules $|m\rangle$; the other field $\Omega_{2}^{\prime}$, in turn, is responsible for associating pairs composed of an atom and a triatomic molecule $|m\rangle$ into tetraatomic molecules $|t\rangle$. For the applied lasers, their frequencies are expressed using the single- and two-photon detunings $\delta$ and $\Delta$, respectively. We remark that the formation of triatomic molecules via the photoassociation of free atoms has been recently reported Schnabel et al. (2021), whereas the formation of tetraatomic molecules through the coupling of an atom-triatomic molecule pair has also been investigated Jing and Jiang (2008). Our aim is to find an exact solution for controlling fast and efficient conversion from ultracold atoms to tetraatomic molecules by generalizing existing N-STIREP protocols.
@@
where $\hat{\psi}_{k}^{\dagger}(\mathbf{r})$ and $\hat{\psi}_{k}^{\dagger}(\mathbf{r})$ are, respectively, the bosonic creation and annihilation operators for state $|k\rangle(k\!=\!a,m,t)$. The indices $i,j=a,m,t$ represent the atomic, triatomic molecular, and tetraatomic molecular states, respectively. By adopting the mean-field approach $\hat{\psi}_{i}\rightarrow\sqrt{n}{\psi}_{i}$ ($n$ denotes the initial atomic density), the Heisenberg motional equation can be described as
$\begin{aligned} i\dot{\psi}_{a}\!&=\!-2K_{a}\psi_{a}\!-\!3\Omega_{1}{\psi_{m}\psi^{\ast}_{a}}^{2}\!+\!\Omega_{2}\psi_{t}{\psi^{*}_{m}},\\ i\dot{\psi}_{m}\!&=\!-2K_{m}\psi_{m}\!-\!(\delta\!+\!i\gamma_{b})\psi_{m}\!-\!\Omega_{1}\psi_{a}^{3}\!+\!\Omega_{2}\psi_{t}{\psi^{\ast}_{a}},\\ i\dot{\psi}_{t}\!&=\!-2K_{t}\psi_{t}\!-\!(\Delta\!+\!\delta\!+\!i\gamma_{t})\psi_{t}\!+\!\Omega_{2}\psi_{m}\psi_{a}.\end{aligned}$ (2)
In the above equations, $\chi_{ij}\!=\!n\chi_{ij}^{\prime},\Omega_{1}\!=\!n\Omega_{1}^{\prime},$ and $\Omega_{2}\!=\!\sqrt{n}\Omega_{2}^{\prime}$ denote the renormalized quantities. $\gamma_{b}$ and $\gamma_{c}$ are introduced to phenomenologically describe the losses of the corresponding states to other undetected states. Nevertheless, we temporarily ignore these losses in the theoretical analysis to clarify the underlying physics. In this case, we have $|\psi_{\mathrm{a}}|^{2}\!+\!3|\psi_{\mathrm{m}}|^{2}\!+\!4|\psi_{\mathrm{t}}|^{2}\!=\!1$. The third-order nonlinearity term $K_{i}\!=\!\sum_{j}\chi_{ij}\left|\psi_{j}\right|^{2}$ characterizes elastic collisions between particles, which gives rise to the dynamical instability effect.
In principle, the effects of third-order nonlinearity can be eliminated by employing resonance locked inverse engineering Dorier et al. (2017); Cao et al. (2023). To this end, a general parametrization with the phase $\alpha$ should be introduced, i.e., $\psi_{\mathrm{a}}\!\mapsto\!\psi_{\mathrm{a}}e^{i\alpha},\psi_{\mathrm{m}}\!\mapsto\!\psi_{\mathrm{m}}e^{\mathrm{3}i\alpha},\psi_{\mathrm{t}}\!\mapsto\!\psi_{\mathrm{t}}e^{\mathrm{4}i\alpha}$. From this, the condition for the resonance locking can be derived as
@@
$\begin{aligned} \Omega_{1}&\!=\!-\frac{\sqrt{3}}{12}\left(\frac{\sin^{2}\theta\!+\!4\cos^{2}\theta}{\cos^{3}\theta\cos^{2}\phi}\dot{\phi}\!+\!\frac{10\tan\theta}{\sin 2\phi\cos\theta}\dot{\theta}\right),\\ \Omega_{2}&\!=\!\frac{\sqrt{3}}{2}\left(\frac{\dot{\theta}}{\sin\phi}\!-\!\frac{\dot{\phi}\tan\theta}{\cos\phi}\right).\end{aligned}$ (6)
To ensure the desired atom-to-tetraatomic molecular conversion, we apply boundary conditions: $0\!\leftarrow\!\phi\!\rightarrow\!0$ and $0\!\leftarrow\!\theta\!\rightarrow\!\pi/2$, where the arrows to the left and right signify the limits when $t\!\rightarrow\!t_{i}$ and $t\!\rightarrow\!t_{f}$, respectively. The controlled parameters $\theta$ and $\phi$ are chosen as
@@ II.2 Result and Discussion
Figure. 2(a) shows the time-dependent evolution of Rabi frequencies used for the conversion of atoms to tetraatomic molecules. In Fig. 2(b), we present the corresponding population dynamics. As anticipated, we achieved an almost perfect conversion efficiency, while maintaining a small transient population in the intermediate state during the interaction. Clearly, our method has the potential to maintain the phase-space density of ultracold mixtures, as a high conversion efficiency is essential for this maintenance.
@@
Figure 3: (Color online) (a) Rabi frequencies as functions of $t$ under the generalized STIREP and the STIRAP. The scaling parameter $\chi=8.4437$ normalizes the two fields to the same magnitude; (b) Conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$ under the (upper surface) generalized STIREP and (lower surface) STIRAP protocols; (c) Conversion efficiency as a function of $T$ under the generalized STIREP; the loss rates are set as $\gamma_{b}=10^{-1}/T$ and $\gamma_{c}=10^{-4}/T$.
@@
Figure. 3(a) displays the time sequence of the Rabi frequencies corresponding to our protocol and STIRAP. It is obvious that, like the Gaussian waveform, the fields used in our protocol are still sufficiently smooth. We plot in Fig. 3(b) the conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$. Here, the upper surface corresponds to our protocol, whereas the lower surface corresponds to STIRAP. As it is observed, our protocol always performs better than STIRAP throughout the whole range of error parameters. In addition, we stress that our protocol allows for faster control over the conversion process. We demonstrate this fact by plotting a curve of the conversion efficiency of our protocol as a function of the evolution time $T$, as depicted in Fig. 3(c). As it turns out, the higher conversion efficiency can be achieved by shortening the evolution time. This phenomenon aligns with the core spirit of STA.
@@
The above calculations demonstrate that the robustness and fast conversion of our protocol facilitate the preparation of ultracold tetraatomic molecules. Nevertheless, it should be stressed that the resulting molecules are usually energetically unstable Kuznetsova et al. (2009), thereby making it highly necessary from a practical standpoint to rapidly transfer them to deeply-bound vibrational states. To this end, we suggest a five-state chainwise-STIREP technique, whose details will be further developed in the following section.

### v3

# Robust, fast, and efficient formation of stable tetratomic molecules from ultracold atoms via generalized stimulated Raman exact passage

@@ Abstract
The study of the conversion of ultracold atoms into molecules has long remained a hot topic in atomic, molecular, and optical physics. However, most prior research has focused on diatomic molecules, with relatively scarce exploration of polyatomic molecules. Here we propose a two-step strategy for the formation of stable ultracold tetratomic molecules. We first suggest a generalized nonlinear stimulated Raman exact passage (STIREP) technique for the coherent conversion of ultracold atoms to tetratomic molecules, which is subsequently followed by a chainwise-STIREP technique to transfer the resulting molecules into a sufficiently stable ground state. Through systematic numerical analysis, we demonstrate that the proposed two-step strategy holds great potential for the robust, fast, and efficient formation of stable ultracold tetratomic molecules.
@@ I Introduction
The successful creation of various ultracold diatomic molecules has stimulated growing interest in polyatomic molecules (UPMs) Rittenhouse and Sadeghpour (2010); Zeppenfeld et al. (2012); Kozyryev et al. (2017); Yang et al. (2022); Vilas et al. (2024); Chen et al. (2024). In contrast to diatomic molecules, UPMs offer additional complexity Carr et al. (2009); Küpper et al. (2009); Patterson et al. (2010), lending themselves to numerous interesting studies in, for instance, cold chemistry Karman et al. (2024), precision measurements Hutzler (2020) and the realization of exotic quantum phases Schmidt et al. (2022), to name a few. In order to tap their full potential, however, long-lived ensembles and full quantum state control serve as prerequisites Langen et al. (2024). Thus, the efficient conversion of ultracold atoms into stable UPMs constitutes an essential requirement, which hitherto remains challenging.
@@
In the ongoing experiments, ultracold molecules are first prepared by association of pre-cooled atoms Chin et al. (2010), either via photo- or the magneto-association Dulieu and Gabbanini (2009); Chandler (2010); Quéméner and Julienne (2012); Zhao and Pan (2022). Subsequently, the formed molecules should be immediately transferred to absolute ground states, it is commonly achieved via the well-known Stimulated Raman adiabatic passage (STIRAP) Drummond et al. (2002); Lang et al. (2008) and its variants, such as multi-state STIRAP Danzl et al. (2010); Mark et al. (2009); Kuznetsova et al. (2008); Jing et al. (2008); Qian et al. (2010); Devolder et al. (2021). In principle, by manipulating the system to evolve along the adiabatic dark state, an efficient and robust state transfer can be obtained Bergmann et al. (1998); Král et al. (2007); Vitanov et al. (2017). Until now, STIRAP has been demonstrated successfully Bergmann et al. (2019), however, its application has been limited to the realm of diatomic molecules Winkler et al. (2007); Lang et al. (2008); Danzl et al. (2008); Aikawa et al. (2010); Ospelkaus et al. (2008); Feng et al. (2009); Liu et al. (2019a); Duda et al. (2023); Zamarski et al. (2025). Directly applying it to UPMs preparation remains challenging. To address this problem, researchers have developed STIRAP variants, for instance, Feshbach/Efimov resonance-assisted STIRAP, to support theoretical investigations into UPMs preparation Jing et al. (2007); Meng et al. (2009); Li and Peng (2011); Dou et al. (2012). Nevertheless, the successful application of such techniques is limited by the adiabatic criterion Zhou et al. (2010), which renders the systems vulnerable to environmental factors, e.g., collisions or decay to outside states.
@@
In order to relax limitations such as adiabaticity related to the STIRAP-based scheme, a technique which is called shortcut-to-adiabaticity (STA) was proposed Chen et al. (2010). By leveraging the STA, the adiabatic process can be expedited while ensuring that the outcomes remain consistent. One particularly successful method for designing such fast routes is counterdiabatic driving Demirplak and Rice (2005, 2008); Berry (2009); Finžgar et al. (2025); Morawetz and Polkovnikov (2025); Dengis et al. (2025a, b), which allows one to construct a modification of an original Hamiltonian to eliminate unwanted nonadiabatic couplings Guéry-Odelin et al. (2019). Currently, STA is widely used in theory to realize fast and robust ultracold molecule formation Zhu and Chen (2021); Zhang and Dou (2021); Cao et al. (2023) and control the selective vibrational population transfer of molecules Masuda and Rice (2015a, b). In a recent study, Zhu etet alalZhu and Chen (2021) proposed a fast-forward scaling-assisted STIRAP method to study the conversion from atoms to diatomic molecules. In another study, Vitanov developed a shortcut multi-state STIRAP technique Vitanov (2020), which may provide an alternative ways for forming stable ground-state ultracold molecules. Compared to the traditional three-level system, the employment of multi-level ones exhibits unique advantages; to illustrate, this strategy proves instrumental in resolving the challenge of weak Franck-Condon factors (FCFs) during the process of both atom-to-molecule conversion Jaksch et al. (2002); Qian et al. (2010) and ground-state molecular transfer Danzl et al. (2010); Kuznetsova et al. (2008). Despite these advances, the major problem is that most prior studies usually require introducing additional couplings for eliminating non-adiabatic effects, which may pose additional experimental challenges. On the other hand, in 2017, Dorier etet alalDorier et al. (2017) suggested an alternative method, namely nonlinear stimulated Raman exact passage (STIREP), for the efficient atom-to-molecule conversion. This method conforms to the spirit of STA, because the control fields are determined by the specific dynamics selected to achieve exact state transfer Laforgue et al. (2022); Dridi et al. (2024). This trick has been extended to the formation of triatomic molecules Cao (2024). Currently, an important question in this realm is whether it is possible to extend this successful tool to form larger polyatomic molecular complexes Dou et al. (2013); Shammout et al. (2025); Quéméner et al. (2023); Cao et al. (2024); Gacesa et al. (2021); Gu et al. (2025), and generalize it to transfer of molecules into deeply-bound states Koch and Moszyński (2008); Danzl et al. (2010); Kuznetsova et al. (2008); de Lima (2024); Mukherjee and Tomza (2025).
@@
The purpose of this paper is to develop a theoretical formalism that extends the existing STIREP protocols to the case that permits control of conversion from atoms to stable tetratomic molecules. We first report a generalization of the STIREP technique for controlling the conversion from ultracold atoms to tetratomic molecules. To transfer the resulting molecules to a sufficiently stable ground state in a short time, we further develop a chainwise-STIREP (abbreviated as C-STIREP hereafter) technique. We present the complete explanations of the underlying mechanism and the corresponding numerical analyses. The results demonstrate that our proposed two-step strategy performs well in both the preparation of ultracold tetratomic molecules and the ground-state molecular transfer.
@@
II COHERENT CONVERSION FROM ATOMS TO TETRATOMIC MOLECULES VIA GENERALIZED NONLINEAR STIREP
@@ II.1 Model and Method
As outlined in the introduction, our top priority is to prepare homonuclear tetratomic molecules. Figure. 1(a) shows a visualization of the process.
Figure 1: (Color online) Schematic showing the coherent conversion from ultracold atoms to tetratomic molecules.
Starting from a Bose condensate of atoms, we apply the field $\Omega_{1}^{\prime}$ to associate free atoms $|a\rangle$ into triatomic molecules $|m\rangle$; the other field $\Omega_{2}^{\prime}$, in turn, is responsible for associating pairs composed of an atom and a triatomic molecule $|m\rangle$ into tetratomic molecules $|t\rangle$. For the applied lasers, their frequencies are expressed using the single-photon and two-photon detunings $\delta$ and $\Delta$, respectively. We remark that the formation of triatomic molecules via the photoassociation of free atoms has been recently reported Schnabel et al. (2021), whereas the formation of tetratomic molecules through the coupling of an atom-triatomic molecule pair has also been investigated Jing and Jiang (2008). Our aim is to find an exact solution for controlling fast and efficient conversion from ultracold atoms to tetratomic molecules by generalizing existing N-STIREP protocols.
@@
where $\hat{\psi}_{k}^{\dagger}(\mathbf{r})$ and $\hat{\psi}_{k}^{\dagger}(\mathbf{r})$ are, respectively, the bosonic creation and annihilation operators for state $|k\rangle(k\!=\!a,m,t)$. The indices $i,j=a,m,t$ represent the atomic, triatomic molecular, and tetratomic molecular states, respectively. By adopting the mean-field approach $\hat{\psi}_{i}\rightarrow\sqrt{n}{\psi}_{i}$ ($n$ denotes the initial atomic density) Ling et al. (2004), the Heisenberg motional equation can be described as
$\begin{aligned} i\dot{\psi}_{a}\!&=\!-2K_{a}\psi_{a}\!-\!3\Omega_{1}{\psi_{m}\psi^{\ast}_{a}}^{2}\!+\!\Omega_{2}\psi_{t}{\psi^{*}_{m}},\\ i\dot{\psi}_{m}\!&=\!-2K_{m}\psi_{m}\!-\!(\delta\!+\!i\gamma_{m})\psi_{m}\!-\!\Omega_{1}\psi_{a}^{3}\!+\!\Omega_{2}\psi_{t}{\psi^{\ast}_{a}},\\ i\dot{\psi}_{t}\!&=\!-2K_{t}\psi_{t}\!-\!(\Delta\!+\!\delta\!+\!i\gamma_{t})\psi_{t}\!+\!\Omega_{2}\psi_{m}\psi_{a}.\end{aligned}$ (2)
In the above equations, $\chi_{ij}\!=\!n\chi_{ij}^{\prime},\Omega_{1}\!=\!n\Omega_{1}^{\prime},$ and $\Omega_{2}\!=\!\sqrt{n}\Omega_{2}^{\prime}$ denote the renormalized quantities. $\gamma_{m}$ and $\gamma_{t}$ are introduced to phenomenologically describe the losses of the corresponding states to other undetected states. Nevertheless, we temporarily ignore these losses in the theoretical analysis to clarify the underlying physics. In this case, we have $|\psi_{\mathrm{a}}|^{2}\!+\!3|\psi_{\mathrm{m}}|^{2}\!+\!4|\psi_{\mathrm{t}}|^{2}\!=\!1$. The third-order nonlinearity term $K_{i}\!=\!\sum_{j}\chi_{ij}\left|\psi_{j}\right|^{2}$ characterizes elastic collisions between particles, which gives rise to the dynamical instability effect Itin and Watanabe (2007); Guérin et al. (2013).
In principle, the effects of third-order nonlinearity can be eliminated by employing resonance locked inverse engineering Dorier et al. (2017); Zhu et al. (2023); Cao et al. (2023). To this end, a general parametrization with the phase $\alpha$ should be introduced, i.e., $\psi_{\mathrm{a}}\!\mapsto\!\psi_{\mathrm{a}}e^{i\alpha},\psi_{\mathrm{m}}\!\mapsto\!\psi_{\mathrm{m}}e^{\mathrm{3}i\alpha},\psi_{\mathrm{t}}\!\mapsto\!\psi_{\mathrm{t}}e^{\mathrm{4}i\alpha}$. From this, the condition for the resonance locking can be derived as
@@
$\begin{aligned} \Omega_{1}&\!=\!-\frac{\sqrt{3}}{12}\left(\frac{1\!+\!3\cos^{2}\theta}{\cos^{3}\theta\cos^{2}\phi}\dot{\phi}\!+\!\frac{10\tan\theta\dot{\theta}}{\sin 2\phi\cos\theta}\right),\\ \Omega_{2}&\!=\!\frac{\sqrt{3}}{2}\left(\frac{\dot{\theta}}{\sin\phi}\!-\!\frac{\dot{\phi}\tan\theta}{\cos\phi}\right).\end{aligned}$ (6)
To ensure the desired atom-to-tetratomic molecular conversion, we apply boundary conditions: $0\!\leftarrow\!\phi\!\rightarrow\!0$ and $0\!\leftarrow\!\theta\!\rightarrow\!\pi/2$, where the arrows to the left and right signify the limits when $t\!\rightarrow\!t_{i}$ and $t\!\rightarrow\!t_{f}$, respectively. The controlled parameters $\theta$ and $\phi$ are chosen as
@@ II.2 Result and Discussion
Figure. 2(a) shows the time-dependent evolution of Rabi frequencies used for the conversion of atoms to tetratomic molecules. In Fig. 2(b), we present the corresponding population dynamics. As anticipated, we achieved an almost perfect conversion efficiency, while maintaining a small transient population in the intermediate state during the interaction. Clearly, our method has the potential to maintain the phase-space density of ultracold mixtures, as a high conversion efficiency is essential for this maintenance.
@@
Figure 3: (Color online) (a) Rabi frequencies as functions of $t$ under the generalized STIREP and the STIRAP. $\chi=8.4437$; (b) Conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$ under the (upper surface) generalized STIREP and (lower surface) STIRAP protocols; (c) Conversion efficiency as a function of $T$ under the generalized STIREP; Parameters: $\gamma_{m}=10^{-1}/T$, $\gamma_{t}=10^{-4}/T$.
@@
Figure. 3(a) displays the time sequence of the Rabi frequencies corresponding to our protocol and STIRAP. It is obvious that, like the Gaussian waveform, the fields used in our protocol are still sufficiently smooth. We plot in Fig. 3(b) the conversion efficiency as a function of $\eta_{1}$ and $\eta_{2}$. Here, the upper surface corresponds to our protocol, whereas the lower surface corresponds to STIRAP. As it is observed, our protocol always performs better than STIRAP throughout the whole range of error parameters. In addition, we stress that our protocol allows for faster control over the conversion process. We demonstrate this fact by plotting a curve of the conversion efficiency of our protocol as a function of the evolution time $T$, as depicted in Fig. 3(c). As it turns out, the higher conversion efficiency can be achieved by shortening the evolution time. This intriguing phenomenon aligns with the core spirit of STA.
@@
The above calculations demonstrate that the robustness and fast conversion of our protocol facilitate the preparation of ultracold tetratomic molecules. Nevertheless, it should be stressed that the resulting molecules are usually energetically unstable Kuznetsova et al. (2009), thereby making it highly necessary from a practical standpoint to rapidly transfer them to deeply-bound vibrational states. To this end, we suggest a five-state C-STIREP technique, whose details will be further analyzed and presented in the following section.